OpenStudy (anonymous):

Jane has a 5 × 7-inch photograph that she wants enlarged to a 10 × 14-inch photograph. What is the scale factor from the original photograph to the enlarged photograph?

Directrix (directrix):

5/10 = 7/14 = (reduces to) ? That is your scale factor from original to enlarged.